OpenStudy (anonymous):

Write an algebra equation for each purchase. I bought an ice cream treat and a bottle of water for $10. Pam purchased an ice cream treat and two bottles of water for $14. Use your equations and reasoning skills to find the price of a bottle of water and an ice cream treat.

OpenStudy (anonymous):

\[x + y = 10\] \[x + 2y = 14\] solve it simultaneously

OpenStudy (anonymous):

do i multiply them/?

OpenStudy (anonymous):

wht do u mean? multiply wt

OpenStudy (anonymous):

What do i do with the numbers 10 and 14

OpenStudy (anonymous):

Can u do the whole problem and show explanation?

OpenStudy (anonymous):

its like this \[x + y = 10\] \[x + 2y = 14\] x is common in both so u have to subtract the second one from 1st x-x + y-2y = 10-14 -y = -4 y=4 which means bottle of water (y) = 4 and then substitute 4 in equation (x+y=10 or x+2y=10) x + 4 = 10 x=6

OpenStudy (anonymous):

Oh ok ;p i get what u did so x=6 and y=4 that would be my answer for the question correct?

OpenStudy (anonymous):

yes it is in this case, bottle cost 4 dollars and icecream cost 6 dollars so when the person buy an ice cream treat and bottle = 6+4 = 10 2nd time he buys an ice cream treat and 2 bottle = 6+ 2(4) = 14 there u get t

OpenStudy (anonymous):

ohh i andustand now thanks alot and gud explanation ;)

OpenStudy (anonymous):

gud explaining are u a teacher?

OpenStudy (anonymous):

nope im a student :) grade 10

OpenStudy (anonymous):

oh kool same i am a sophmore

OpenStudy (anonymous):

:)